OpenStudy (anonymous):

MEDAL AND FAN!!!!!! Evaluate the function rule for the given value. It's in the comments

OpenStudy (anonymous):

\[y=15.3^{x} for x=-3\]

OpenStudy (anonymous):

The multiple choices are A) 5/27 B) 5/3 C) 5/9 D) -135

OpenStudy (ineedhelplz):

The answer is B :)

jhonyy9 (jhonyy9):

for example x^(-1) = 1/x hope this will help you

OpenStudy (anonymous):

thx ineedhelplz

OpenStudy (ineedhelplz):

or that

jhonyy9 (jhonyy9):

@Ineedhelplz but why dont learn the way how you can calculi the right answer please ?

jhonyy9 (jhonyy9):

why dont teach ???? - sorry

OpenStudy (ineedhelplz):

I got the answer wrong :(

OpenStudy (ineedhelplz):

It's hard to teach :/

jhonyy9 (jhonyy9):

@beckcyla do you need again help for this exercise please ?

jhonyy9 (jhonyy9):

@Ineedhelplz sorry but not is hard to teach than you know the way how you make it for the asker can understanding the right way

jhonyy9 (jhonyy9):

15,3^(-3) = ? than you use formula what i have wrote to you above